# Review: Quantum Metrology and Sensing with Many-Body Systems

### v2

@@ IV Equilibrium Quantum Sensing: Second Order Quantum Phase Transition
where $\eta$ is the scaling exponent of the long-range correlation function of the order parameter, as defined by $\langle\sigma_{i}^{x}\sigma_{i+r}^{x}\rangle\propto r^{2-d-\eta}$, at critical point for a $d$-dimensional model, and $d_{m}=\min(d,d_{c})$, with upper critical dimension $d_{c}=3$ for the transverse field Ising chain. The results, plotted in Fig. 3, attest that sub-shot-noise scaling is indeed achievable with such probes. Similar results have also been found in Ref. Li et al. (2024) in which interferometric quantum enhanced sensitivity, using the critical ground state of a many-body system, can be exploited for detecting quantum phase transitions.
@@
Figure 3: Sub-shot-noise scaling with critical quantum transverse field Ising chain probes for various dimensions. In (b) and (c) $x$-axis is $N_{A}$, the number of spins inside any subsystem $A$ as depicted in (a); $y$-axis is (b) variance of the estimator of $y$-component of collective spin operator $J=\sum_{i}S_{i}$ normalized by $N_{A}$, and (c) collective spin squeezing parameter $\xi_{R}^{2}$. Figure taken from Ref. Frérot and Roscilde (2018).
@@
$I_{Q}(h){=}N^{\alpha/\nu}g\left(N^{1/\nu}(\theta-\theta_{c})\right),$ (41)
where $g(\cdot)$ is an arbitrary function. This ansatz is routinely used for determining critical exponents $\alpha$ and $\nu$ using finite size systems. Plotting the rescaled QFI, namely $N^{-\alpha/\nu}I_{Q}(h)$, versus $N^{1/\nu}(h-h_{c})$ collapses all the curves for different system sizes for the right choice of critical parameters, namely $(\theta_{c},\alpha,\nu)$. Since the two ansatzes in Eq. (40) and Eq. (41) describe the same quantity, they are not independent. In fact, by factorizing $N^{\beta}$ form Eq. (40), one can show that the identicality of both ansatzes imposes the following identity between the critical exponents
@@
This shows that the three critical exponents are not independent from each other and knowing two of them is enough for determining the other. While the identity in Eq. (42) is valid for one-dimensional systems, its generalization to $d$ dimensional probes is obtained as $\beta=\frac{\alpha}{d\nu}$ Schwandt et al. (2009); Albuquerque et al. (2010); Rams et al. (2018). This is a remarkable discovery which shows that unlike the interferometry-based sensing scenarios, e.g. using GHZ states, the precision of criticality-based quantum sensing is not bounded by the Heisenberg scaling. In fact, super-Heisenberg scaling with $\beta>2$ can be achieved if $\alpha>2d\nu$. It is worth mentioning that for all realization of the second-order phase transition in the ground state, one can show that $\alpha{=}2$ Schwandt et al. (2009); Albuquerque et al. (2010). However, some non-ground state continous phase transitions behave like the second-order phase transition with $\alpha$ different from 2 He et al. (2023). Therefore, for keeping generality of the paper, we keep $\alpha$ as a general parameter in our notations.
@@ IV.1 Free fermionic many-body probes
The above discussion on second-order quantum phase transitions is very general. It will be more insightful to go through some examples and explicitly determine the scaling of QFI, namely $I_{Q}\sim N^{\beta}$, at the vicinity of the critical point. Free fermionic systems are an important class of many-body systems in condensed matter physics. In particular, they are analytically solvable which makes them an excellent toy model for understanding several complex phenomena, such as criticality. In addition, many spin chain systems can also be mapped to free fermion systems through Jordan-Wigner transformation and thus become solvable Sachdev and Safari (2011). A general free fermion model with $N$ particles can be described by Hamiltonian $\hat{H}(\theta)$ which in the Nambu formalism notation is written as (Mbeng et al., 2020)
$\hat{H}(\theta)=\Psi^{\dagger}\tilde{H}(\theta)\Psi,$ (43)
where $\Psi^{\dagger}=\left(\hat{c_{1}}^{\dagger}\hat{c_{1}}^{\dagger}\dots\hat{c_{N}}^{\dagger}\hat{c_{1}}\hat{c_{2}}\dots\hat{c_{N}}\right)$ is a $2N\times 1$ vector with $\hat{c_{1}}(\hat{c_{1}}^{\dagger})$ being bare-particle annihilation (creation) operators and $\tilde{H}(\theta)$ being the $2N\times 2N$ Hermitian matrix containing information about $\theta$. The collection of normalized eigenvectors $\{|E_{k}\rangle\}$ of $\tilde{H}$ are expressible in terms of $N\times N$-matrices $U$ and $V$ as
@@
This closed form of QFI can be used to study the scaling of QFI in free fermion systems. Another useful perspective for calculation of QFI for systems governed by generic quadratic fermionic Hamiltonians was provided by Carollo et al (Carollo et al., 2018a), where the SLD of fermionic Gaussian states associated with such quadratic Hamiltonians was considered. In this case, instead of the Dirac-representation of fermions $\{c_{i},c_{j}^{\dagger}\}=\delta_{ij},\{c_{i},c_{j}\}=0$, it is more convenient to work with the Hermitian Majorana representation given by fermion operators $\omega$ such that $\omega_{2j-1}=c_{j}+c_{j}^{\dagger},\omega_{2j}=i(c_{j}-c_{j}^{\dagger})$. Fermionic Gaussian states with $n$-fermionic modes and parametrized by the sensing parameter $\theta$, can then be defined as states of the form $\rho(\theta)=\exp(-\frac{i}{4}\omega^{T}\Omega(\theta)\omega)/\operatorname{Tr}\left[\exp(-\frac{i}{4}\omega^{T}\Omega(\theta)\omega)\right]$, where $\Omega(\theta)$ is a $2n\times 2n$ real anti-symmetric matrix and $\omega=(\omega_{1},\omega_{2},...,\omega_{2n})$ are the $2n$ Majorana fermions. Owing to the above Gaussian structure, such states can be completely described in terms of the second moment, i.e., the two-point correlation matrix $\Gamma_{jk}(\theta)=\frac{1}{2}\operatorname{Tr}[\rho(\theta)(\omega_{j}\omega_{k}{-}\omega_{k}\omega_{j})]{=}\tanh{(i\Omega_{jk}(\theta)/2)}$. Following the established mathematical approach for photonic Gaussian states (Serafini, 2017), Carollo et al used the ansatz that the SLD in this case is also given by the quadratic form
$L(\theta)=\frac{1}{2}\omega^{T}K(\theta)\omega+\xi^{T}(\theta)\omega+\eta(\theta),$ (46)
@@ IV.2 Bosonic many-body probes
$H=-t\sum_{i}\left(b_{i}^{\dagger}b_{i+1}+b_{i+1}^{\dagger}b_{i}\right)-\mu\sum_{i}n_{i}+\frac{U}{2}\sum_{i}n_{i}\left(n_{i}-1\right)$ (48)
where $b_{i}(b_{i}^{\dagger})$-s are the Bosonic annihilation (creation) operators, $n_{i}{=}b_{i}^{\dagger}b_{i}$ is the number operator at site $i$, and $t,U,\mu$ are the hopping, onsite interaction, and chemical potential paremeters, respectively. The Bose-Hubbard model can be implemented by loading cold bosonic atoms in Harmonic oscillator traps (Jaksch et al., 1998; Jaksch and Zoller, 2005; Lewenstein et al., 2007). At $T=0$, the phase diagram consists of a Mott insulating phase ($t\ll U$) and a Superfluid phase ($t\gg U$). The quantum phase transition between these phases from the Mott insulator phase to the Superfluid phase involves the spontaneous breaking of the $U(1)$ symmetry of the Bose-Hubbard Hamiltonian. Quantum sensing in these atomic systems combines the theoretical formalism of photon-interfermometry based metrology with reasonable ease of producing quantum states and robustness against particle loss. Considering a single-mode description of the resulting interacting BEC, where atom-atom interactions are described by a non-linear correction to the Harmonic trap potential, Gross et al obtained the first experimental demonstration of quantum-enhanced metrology in cold Bosonic atoms Gross et al. (2010). Dunningham et al proposed the use of number-correlated squeezed BEC states (Dunningham et al., 2002), which are more robust to particle loss than NOON states. Another scheme by Dunningham and Burnett suggested that one can recreate the action of a (multiport) beam-splitter in an interferometer by simply tweaking the potential barrier between sites, thus enabling quantum-enhanced sensitivity (Dunningham and Burnett, 2004; Cooper et al., 2009). Heisenberg-limited quantum analogs of inertial guidance systems, also known as atomic gyroscopes, have been proposed in optical ring lattices (Cooper et al., 2010). A multiconfigurational Hartree type self-consistent theoretical method for calculation of quantum metrological quantities for interacting trapped Bosonic systems has also been proposed Baak and Fischer (2024) . Precision estimation of chemical potential has also been studied in Bosonic quantum gases Marzolino and Braun (2013). While strictly not many-body systems, criticalities associated with superradiant phase transitions engineered in optical cavities with Rabi or Jaynes-Cummings type interactions have also received renewed attention for the possibility of quantum-enhanced sensing  Garbe et al. (2020). Moreover, effect of engineering anisotropies between coupling strengths of rotating wave and counter-rotating atom-cavity interactions of such models have also been studied in detail Wang et al. (2016); Zhu et al. (2023).
@@ IV.3 Spin chain probes with criticality-enhanced sensitivity
$H=h\underbrace{\sum_{i=1}^{N}\sigma^{z}_{i}}_{H_{1}}+J\underbrace{\sum_{i=1}^{N-1}\sigma^{x}_{i}\sigma^{x}_{i+1}}_{H_{2}}.$ (49)
@@
By factoring $J$ and defining $\theta{=}h/J$, the Hamiltonian takes the general form of Eq. (36). One can analytically solve the transverse Ising model Eq. (49) through Jordan-Wigner transformation (Pfeuty, 1970) which connects spin operators to fermionic ones as $c_{i}=\prod_{j=1}^{i-1}(-\sigma^{z}_{i})\sigma^{-}_{i}$ (and equivalently $c_{i}^{\dagger}=\prod_{j=1}^{i-1}(-\sigma^{z}_{i})\sigma^{+}_{i}$), where $\sigma^{\pm}{=}(\sigma^{x}\pm i\sigma^{y})/2$. The corresponding QFI for the dimensionless parameter $\theta$ has been analytically calculated for the ground state (Zanardi et al., 2008; Invernizzi et al., 2008) and has been shown to scale quadratically with system size, namely $I_{Q}\sim N^{2}$ near the critical point $h/J\rightarrow\pm 1$. Recently, a modular approach effectively augmenting these systems with periodic coupling defects (Mukhopadhyay and Bayat, 2024) has been proposed, which exploits the fact that multiple criticalities can be created in such a construction, all of which inherit the critical scaling behavior of the original chain. Alternatively, even for uniform chains, extending the ensemble measurement setup to real-time feedback control, i.e., adaptive strategies, allows one to overcome shot-noise scaling across the phase diagram  Salvia et al. (2023).
@@
Other quantum spin models have also been extensively studied in the literature for determination of quantum-enhanced scaling. In addition to their relevance for quantum sensing, such QFI-based analysis often reveals the nature of multipartite entanglement in many-body quantum systems. These include the 1D quantum XY models (Liu et al., 2013; Skotiniotis et al., 2015; Ye et al., 2016) with Dzyaloshinsky-Moriya interactions (Ozaydin and Altintas, 2015; Ye et al., 2018; Ozaydin and Altintas, 2020), XXZ models (Liu et al., 2015) which show Kosterlitz-Thouless phase transition (Zheng et al., 2015), Heisenberg-type rotationally symmetric models (Ye et al., 2020; Lambert and Sørensen, 2019), and even applied to the quantum Hall probes Giraud et al. (2024). See Ref. (Gu, 2010) for more references and earlier reviews. In addition to spinless fermion models, metrological advantage was also investigated for short-range spinful Fermi gases and QFI was shown to be an order parameter in such systems (Lucchesi and Chiofalo, 2019). Higher dimensional spin chains have also been shown to improve magnetometry with arbitrary spin-number transverse Ising models Singh et al. (2024). Effect of lattice imperfections and disorder was considered in Ref. (Wei, 2019), whereby QFI analysis reveals that the two paradigmatic quantum disordered 1D lattices, namely Anderson  (Anderson, 1958) and Aubry-Andre models (Aubry and André, 1980) belong to different universality classes. Experimental demonstration of quantum-enhanced critical sensing has also been recently performed (Liu et al., 2021) with a small NMR system. Extensions to multiparameter estimation by calculating the QFI matrix have also been considered (Bakmou et al., 2019; Di Fresco et al., 2022) along with experimental demonstration for NMR systems (Jiang et al., 2021c). Departure from $T=0$, i.e., ground state assumption has been considered in several directions. One direction is to consider metrology in finite-temperatures. Similar to other markers of quantumness, presence of a quantum critical region fanning out from the critical point at $T=0$ has been detected with QFI (Gabbrielli et al., 2018). Another popular direction of departure, namely considering driven quantum many-body sensors instead of probes at equilibrium has been discussed in a later section of this review.

### v3

@@ IV Equilibrium Quantum Sensing: Second Order Quantum Phase Transition
where $\eta$ is the scaling exponent of the long-range correlation function of the order parameter, as defined by $\langle\hat{\sigma}_{i}^{x}\hat{\sigma}_{i+r}^{x}\rangle\propto r^{2-d-\eta}$, at critical point for a $d$-dimensional model, and $d_{m}=\min(d,d_{c})$, with upper critical dimension $d_{c}=3$ for the transverse field Ising chain. The results, plotted in Fig. 3, attest that sub-shot-noise scaling is indeed achievable with such probes. Similar results have also been found in Ref. Li et al. (2024) in which interferometric quantum-enhanced sensitivity, using the critical ground state of a many-body system, can be exploited for detecting quantum phase transitions.
@@
Figure 3: Sub-shot-noise scaling with critical quantum transverse field Ising chain probes for various dimensions. In (b) and (c) $x$-axis is $N_{A}$, the number of spins inside any subsystem $A$ as depicted in (a); $y$-axis is (b) variance of the estimator of $y$-component of collective spin operator $\hat{J}^{y}=\sum_{i}\hat{S}^{y}_{i}$ normalized by $N_{A}$, and (c) collective spin squeezing parameter $\xi_{R}^{2}$. Figure taken from Ref. Frérot and Roscilde (2018).
@@
$I_{Q}(\theta){=}N^{\alpha/\nu}g\left(N^{1/\nu}(\theta-\theta_{c})\right),$ (41)
where $g(\cdot)$ is an arbitrary function which depends on the model and the quantity of interest. This ansatz is routinely used for determining critical exponents $\alpha$ and $\nu$ using finite size systems. Plotting the rescaled QFI, namely $N^{-\alpha/\nu}I_{Q}(\theta)$, versus $N^{1/\nu}(\theta-\theta_{c})$ collapses all the curves for different system sizes for the right choice of critical parameters, namely $(\theta_{c},\alpha,\nu)$. Since the two ansatzes in Eq. (40) and Eq. (41) describe the same quantity, they are not independent. In fact, by factorizing $N^{\beta}$ from Eq. (40), one can show that the identicality of both ansatzes imposes the following identity between the critical exponents
@@
This shows that the three critical exponents are not independent from each other and knowing two of them is enough for determining the other. While the identity in Eq. (42) is valid for one-dimensional systems, its generalization to $d$ dimensional probes is obtained as $\beta=\frac{\alpha}{d\nu}$ Schwandt et al. (2009); Albuquerque et al. (2010); Rams et al. (2018). This is a remarkable discovery which shows that unlike the interferometry-based sensing scenarios, e.g. using GHZ states, the precision of criticality-based quantum sensing is not bounded by the Heisenberg scaling. In fact, quantum-enhanced sensitivity with $\beta>2$ can be achieved if $\alpha>2d\nu$. It is worth mentioning that for all realization of the second-order phase transition in the ground state, one can show that $\alpha{=}2$ Schwandt et al. (2009); Albuquerque et al. (2010). However, some non-ground state continuous phase transitions behave like the second-order phase transition with $\alpha$ different from 2 He et al. (2023). Therefore, for keeping generality of the paper, we keep $\alpha$ as a general parameter in our notations.
It is worth emphasizing that in criticality-based quantum sensing, quantum-enhanced precision can only be harnessed in the notion of local sensing, where significant prior information about the unknown parameter is available. Two primary reasons for this are: (i) as discussed above, quantum-enhanced precision can only be achived when the probe is operating near its criticality, namely $\theta{\sim}\theta_{c}$; and (ii) the optimal measurement basis, in general, depends on the unknown parameter $\theta$ and thus prior information is required to perform a measurement which is nearly optimal. In other words, when the parameter $\theta$ is roughly known and the purpose of sensing is to estimate its value with more precision, one can tune the probe to operate near its criticality and then perform a measurement which is optimal based on the prior value of the parameter. In such scenario, one can obtain the exact value of the parameter $\theta$ with achieving quantum-enhanced precision.
@@ IV.1 Free fermionic many-body probes
The above discussion on second-order quantum phase transitions is very general. It will be more insightful to go through some examples and explicitly determine the scaling of QFI, namely $I_{Q}\sim N^{\beta}$, at the vicinity of the critical point. Free fermionic systems are an important class of many-body systems in condensed matter physics. In particular, they are analytically solvable which makes them an excellent toy model for understanding several complex phenomena, such as criticality. In addition, many spin chain systems can also be mapped to free fermion systems through Jordan-Wigner transformation and thus become solvable Sachdev (1999). A general free fermion model with $N$ particles can be described by Hamiltonian $\hat{H}(\theta)$ which in the Nambu formalism notation is written as (Mbeng et al., 2020)
$\hat{H}(\theta)=\hat{\Phi}^{\dagger}\hat{H}^{\prime}(\theta)\hat{\Phi},$ (43)
where $\hat{\Phi}^{\dagger}=\left(\hat{c}_{1}^{\dagger}\hat{c}_{2}^{\dagger}\dots\hat{c}_{N}^{\dagger}\hat{c}_{1}\hat{c}_{2}\dots\hat{c}_{N}\right)$ is a $2N\times 1$ vector with $\hat{c}_{j}(\hat{c}_{j}^{\dagger})$ being bare-particle annihilation (creation) operators and $\hat{H}^{\prime}$ being the $2N\times 2N$ Hermitian matrix containing information about $\theta$. The collection of normalized eigenvectors $\{|E_{k}\rangle\}$ of $\hat{H}^{\prime}$ are expressible in terms of $N\times N$-matrices $U$ and $V$ as
@@
This closed form of QFI can be used to study the scaling of QFI in free fermion systems. Another useful perspective for calculation of QFI for systems governed by generic quadratic fermionic Hamiltonians was provided by Carollo et al (Carollo et al., 2018a), where the SLD of fermionic Gaussian states associated with such quadratic Hamiltonians was considered. In this case, instead of the Dirac-representation of fermions $\{\hat{c}_{i},\hat{c}_{j}^{\dagger}\}=\delta_{ij},\{\hat{c}_{i},\hat{c}_{j}\}=0$, it is more convenient to work with the Hermitian Majorana representation given by fermion operators $\omega$ such that $\hat{\omega}_{2j-1}=\hat{c}_{j}+\hat{c}_{j}^{\dagger}$, and $\hat{\omega}_{2j}=i(\hat{c}_{j}-\hat{c}_{j}^{\dagger})$. Fermionic Gaussian states with $n$-fermionic modes and parametrized by the sensing parameter $\theta$, can then be defined as states of the form $\rho(\theta)=\exp(-\frac{i}{4}\hat{\bm{\omega}}^{T}\Omega(\theta)\hat{\bm{\omega}})/\operatorname{Tr}\left[\exp(-\frac{i}{4}\hat{\bm{\omega}}^{T}\Omega(\theta)\hat{\bm{\omega}})\right]$, where $\Omega(\theta)$ is a $2n\times 2n$ real anti-symmetric matrix and $\hat{\bm{\omega}}=(\hat{\omega}_{1},\hat{\omega}_{2},...,\hat{\omega}_{2n})$ are the $2n$ Majorana fermions. Owing to the above Gaussian structure, such states can be completely described in terms of the second moment, i.e., the two-point correlation matrix $\Gamma_{jk}(\theta)=\frac{1}{2}\operatorname{Tr}[\rho(\theta)(\hat{\omega}_{j}\hat{\omega}_{k}{-}\hat{\omega}_{k}\hat{\omega}_{j})]{=}\tanh{(i\Omega_{jk}(\theta)/2)}$. Following the established mathematical approach for photonic Gaussian states (Serafini, 2017), Carollo et al used the ansatz that the SLD in this case is also given by the quadratic form
$\hat{L}(\theta)=\frac{1}{2}\hat{\bm{\omega}}^{T}K(\theta)\hat{\bm{\omega}}+\xi^{T}(\theta)\hat{\bm{\omega}}+\eta(\theta)$ (46)
@@ IV.2 Bosonic many-body probes
$H=-J\sum_{i}\left(\hat{b}_{i}^{\dagger}\hat{b}_{i+1}+\hat{b}_{i+1}^{\dagger}\hat{b}_{i}\right)-\mu\sum_{i}\hat{n}_{i}+\frac{U}{2}\sum_{i}\hat{n}_{i}\left(\hat{n}_{i}-1\right)$ (48)
where $\hat{b}_{i}(\hat{b}_{i}^{\dagger})$-s are the Bosonic annihilation (creation) operators, $\hat{n}_{i}{=}\hat{b}_{i}^{\dagger}\hat{b}_{i}$ is the number operator at site $i$, and $J,U,\mu$ are the hopping, onsite interaction, and chemical potential paremeters, respectively. The Bose-Hubbard model can be implemented by loading cold bosonic atoms in Harmonic oscillator traps (Jaksch et al., 1998; Jaksch and Zoller, 2005; Lewenstein et al., 2007). At zero temperature, the phase diagram consists of a Mott insulating phase ($J\ll U$) and a Superfluid phase ($J\gg U$). The quantum phase transition between these phases from the Mott insulator phase to the Superfluid phase involves the spontaneous breaking of the $U(1)$ symmetry of the Bose-Hubbard Hamiltonian. Quantum sensing in these atomic systems combines the theoretical formalism of photon-interfermometry based metrology with reasonable ease of producing quantum states and robustness against particle loss. Considering a single-mode description of the resulting interacting BEC, where atom-atom interactions are described by a non-linear correction to the Harmonic trap potential, Gross et al obtained the first experimental demonstration of quantum-enhanced metrology in cold Bosonic atoms Gross et al. (2010). Dunningham et al proposed the use of number-correlated squeezed BEC states (Dunningham et al., 2002), which are more robust to particle loss than $N00N$ states. Another scheme by Dunningham and Burnett suggested that one can recreate the action of a (multiport) beam-splitter in an interferometer by simply tweaking the potential barrier between sites, thus enabling quantum-enhanced sensitivity (Dunningham and Burnett, 2004; Cooper et al., 2009). Heisenberg-limited quantum analogs of inertial guidance systems, also known as atomic gyroscopes, have been proposed in optical ring lattices (Cooper et al., 2010). A multiconfigurational Hartree type self-consistent theoretical method for calculation of quantum metrological quantities for interacting trapped Bosonic systems has also been proposed Baak and Fischer (2024) . Precision estimation of chemical potential has also been studied in Bosonic quantum gases Marzolino and Braun (2013). While strictly not many-body systems, criticalities associated with superradiant phase transitions engineered in optical cavities with Rabi or Jaynes-Cummings type interactions have also received renewed attention for the possibility of quantum-enhanced sensing  Garbe et al. (2020). Moreover, effect of engineering anisotropies between coupling strengths of rotating wave and counter-rotating atom-cavity interactions of such models have also been studied in detail Wang et al. (2016); Zhu et al. (2023).
@@ IV.3 Spin chain probes with criticality-enhanced sensitivity
$\hat{H}=h\underbrace{\sum_{i=1}^{N}\hat{\sigma}^{z}_{i}}_{\hat{H}_{1}}+J\underbrace{\sum_{i=1}^{N-1}\hat{\sigma}^{x}_{i}\hat{\sigma}^{x}_{i+1}}_{\hat{H}_{2}}.$ (49)
@@
By factoring $J$ and defining $\theta{=}h/J$, the Hamiltonian takes the general form of Eq. (36). One can analytically solve the transverse Ising model Eq. (49) through Jordan-Wigner transformation (Pfeuty, 1970) which connects spin operators to fermionic ones as $\hat{c}_{i}=\prod_{j=1}^{i-1}(-\hat{\sigma}^{z}_{i})\hat{\sigma}^{-}_{i}$ (and equivalently $\hat{c}_{i}^{\dagger}=\prod_{j=1}^{i-1}(-\hat{\sigma}^{z}_{i})\hat{\sigma}^{+}_{i}$), where $\hat{\sigma}^{\pm}{=}(\hat{\sigma}^{x}\pm i\hat{\sigma}^{y})/2$. The corresponding QFI for the dimensionless parameter $\theta$ has been analytically calculated for the ground state (Zanardi et al., 2008; Invernizzi et al., 2008) and has been shown to scale quadratically with system size, namely $I_{Q}\sim N^{2}$ near the critical point $h/J\rightarrow\pm 1$. Recently, a modular approach effectively augmenting these systems with periodic coupling defects (Mukhopadhyay and Bayat, 2024) has been proposed, which exploits the fact that multiple criticalities can be created in such a construction, all of which inherit the critical scaling behavior of the original chain. Alternatively, even for uniform chains, extending the ensemble measurement setup to real-time feedback control, i.e., adaptive strategies, allows one to overcome shot-noise scaling across the phase diagram  Salvia et al. (2023).
@@
Other quantum spin models have also been extensively studied in the literature for determination of quantum-enhanced scaling. In addition to their relevance for quantum sensing, such QFI-based analysis often reveals the nature of multipartite entanglement in many-body quantum systems. These include the 1D quantum XY models (Liu et al., 2013; Skotiniotis et al., 2015; Ye et al., 2016) with Dzyaloshinsky-Moriya interactions (Ozaydin and Altintas, 2015; Ye et al., 2018; Ozaydin and Altintas, 2020), XXZ models (Liu et al., 2015) which show Kosterlitz-Thouless phase transition (Zheng et al., 2015), Heisenberg-type rotationally symmetric models (Ye et al., 2020; Lambert and Sørensen, 2019), and even applied to the quantum Hall probes Giraud et al. (2024). See Ref. (Gu, 2010) for more references and earlier reviews. In addition to spinless fermion models, metrological advantage was also investigated for short-range spinful Fermi gases and QFI was shown to be an order parameter in such systems (Lucchesi and Chiofalo, 2019). Higher dimensional spin chains have also been shown to improve magnetometry with arbitrary spin-number transverse Ising models Singh et al. (2024). Effect of lattice imperfections and disorder was considered in Ref. (Wei, 2019), whereby QFI analysis reveals that the two paradigmatic quantum disordered 1D lattices, namely Anderson  (Anderson, 1958) and Aubry-Andre models (Aubry and André, 1980) belong to different universality classes. Experimental demonstration of quantum-enhanced critical sensing has also been recently performed (Liu et al., 2021b) with a small NMR system. Extensions to multiparameter estimation by calculating the QFI matrix have also been considered (Bakmou et al., 2019; Di Fresco et al., 2022) along with experimental demonstration for NMR systems (Jiang et al., 2021c). Departure from $T=0$, i.e., ground state assumption has been considered in several directions. One direction is to consider metrology in finite-temperatures. Similar to other markers of quantumness, presence of a quantum critical region fanning out from the critical point at $T=0$ has been detected with QFI (Gabbrielli et al., 2018). Another popular direction of departure, namely considering driven quantum many-body sensors instead of probes at equilibrium has been discussed in a later section of this review.